# To Err Is Human, but Llamas Can Learn It Too Анотацiя

### v1

@@
This study explores enhancing grammatical error correction (GEC) through artificial error generation (AEG) using language models (LMs). Specifically, we fine-tune Llama 2-based LMs for error generation and find that this approach yields synthetic errors akin to human errors. Next, we train GEC Llama models with the help of these artificial errors and outperform previous state-of-the-art error correction models, with gains ranging between 0.8 and 6 F points across all tested languages (German, Ukrainian, and Estonian). Moreover, we demonstrate that generating errors by fine-tuning smaller sequence-to-sequence models and prompting large commercial LMs (GPT-3.5 and GPT-4) also results in synthetic errors beneficially affecting error generation models.
@@ 1 Introduction
The grammatical error correction (GEC) task aims to correct spelling and grammatical errors in the input text. The best-performing approaches to this task currently use neural networks (Junczys-Dowmunt et al., 2018; Omelianchuk et al., 2020; Rothe et al., 2021, and several others), which are known to be data-hungry. At the same time, openly available human error correction data is severely limited even for high-resource languages like German, Arabic, and Czech Bryant et al. (2023).
@@
The lack of error correction data is commonly addressed through the creation of synthetic data, where errors are automatically added into correct sentences – also called artificial error generation (AEG). The most common approach to AEG is applying random probabilistic perturbation (deletion, insertion, replacement) of words and/or characters in the correct sentence Zhao et al. (2019); Grundkiewicz et al. (2019); Rothe et al. (2021), alternatives include usage of intricate hand-crafted rules and confusion sets Rozovskaya and Roth (2010); Xu et al. (2019); Kara et al. (2023); Bondarenko et al. (2023) and automatically learning to generate errors Xie et al. (2018); Kiyono et al. (2019); Stahlberg and Kumar (2021) – also referred to as back-translation (BT)***by analogy with the machine translation technique Sennrich et al. (2016). However, to the best of our knowledge, none of the related work on AEG makes use of pre-trained foundation models.
@@
This gap is precisely the focus of the present work: using pre-trained language models for synthetic error generation. We approach the task by fine-tuning open language models (LMs) that are based on Llama 2 Touvron et al. (2023) and show that this can result in successful AEG results even when very limited amounts of human error data are available. Our analysis shows that the resulting errors are much more similar to natural human errors. We also compare the approach to prompting commercial LMs (GPT-3.5 and GPT-4: OpenAI, 2023) to perform AEG, as well as include other open models commonly employed for GEC and tune them for AEG: mT5 Rothe et al. (2021); Palma Gomez et al. (2023) and NLLB Luhtaru et al. (2024). The details of our proposed methodology are given in Section 3.
@@
Our final goal and evaluation setting is improving grammatical error correction for low-resource languages. In particular, we focus on German, Ukrainian, and Estonian GEC. For error correction, we also fine-tuned Llama 2 and compared it to the prompting of variants of GPT-4. Our experimental results show that Llama-based language models with fewer learned parameters can sometimes beat state-of-the-art results achieved with a bigger model. When pre-trained on our LM-generated synthetic errors, the resulting GEC models achieve the best current results on the included benchmarks in all three evaluated cases, including previous state-of-the-art and 4-shot GPT-4.
@@
We show that pre-trained language models can be fine-tuned to generate high-quality synthetic errors.
@@ 2 Related Work
The use of synthetic data is a common concept in GEC. The first effective neural method proposed by Junczys-Dowmunt et al. (2018) approaches GEC as low-resource Machine Translation (MT), making it a relatively resource-heavy method encouraging synthetic data generation. Over the years, there have been different approaches to deliberately introducing errors into monolingual text, like rule-based and probabilistic methods, methods based on confusion sets and error patterns, models trained for error generation and using round-trip translation Bryant et al. (2023).
@@
Learned methods of error generation typically require more resources. Before the widespread adoption of transformers and MT, various studies explored alternative approaches for training models for error generation. For instance, Felice and Yuan (2014) and Rei et al. (2017) utilized statistical machine translation to generate errors, while Xie et al. (2018) and Yuan et al. (2019) experimented with convolutional neural networks (CNNs) for this purpose. Additionally, Kasewa et al. (2018) investigated using RNN-based sequence-to-sequence models with attention mechanisms.
Moving towards more modern MT architectures, Htut and Tetreault (2019) tested various model frameworks, including transformers, and Kiyono et al. (2019) specifically employed transformer models. Both of the latter studies trained models from scratch, utilizing datasets ranging from approximately 500,000 to over a million error correction examples to train the artificial error generation system. In contrast, our work generates up to 1 million sentences with synthetic error while using between 9k and 33k human error sentences to fine-tune the base models.
@@
In addition, Oda (2023) shows that generating new targets for synthetic datasets can be beneficial. Fang et al. (2023) argue that translationese can be closer in domain to language learner’s text than traditional monolingual corpora, which can cause domain mismatch problems.
@@ 3 Methodology and Experiments
The primary target of our work is to apply generative language models to artificial error generation (AEG) via fine-tuning. Additionally, we experiment with prompting large language models to perform the same task and include two seq2seq models that are fine-tuned to do the same.
The efficiency of proposed AEG solutions is evaluated using them to improve grammatical error correction (GEC). Thus, we also fine-tune generative LMs to perform the GEC task and compare the results to prompting-based GEC results and related work.
The general pipeline of our approach is straight-forward:
@@
Fine-tune an LM to generate errors using human error data, with correct sentences as input and sentences with errors as output.
@@
Apply that AEG LM to correct sentences in order to add a synthetically erroneous counterpart
@@
Fine-tune an LM on that synthetic dataset to correct grammatical errors. Equivalent to Step 1, with the sentence pair direction reversed.
@@
Continue fine-tuning GEC LM on the (typically smaller) dataset with human errors.
@@ 3.1 Data
We use two distinct types of data in our work. Firstly, we rely on datasets containing examples of grammatical error corrections to train our error generation systems and correction models. Secondly, we incorporate monolingual data to create synthetic datasets by introducing errors. See an overview of used data in Table 1.
@@ 3.2 Models and Training
Llama-2-based models. We fine-tune models that have been enhanced with bilingual capabilities using continued pre-training from Llama-2-7B (Touvron et al., 2023). For Estonian, we use Llammas-base†††huggingface.co/tartuNLP/Llammas-base, and for German, LeoLM‡‡‡huggingface.co/LeoLM/leo-hessianai-7b. For Ukrainian, we apply continued pre-training to replicate the conditions of Estonian LM by training with 5B tokens from CulturaX (Nguyen et al., 2023) with 25% of the documents being in English and the rest in Ukrainian. For GEC and AEG fine-tuning, we formatted the training data with a prompt (see Table 8 and 9) loosely based on Alpaca (Taori et al., 2023). During fine-tuning, the loss is calculated on the tokens of the correct sentence. Fine-tuning details (including hyperparameters) are discussed in Appendix B.1.
Other models we use are NLLB Team et al. (2022) and mT5 Xue et al. (2021). Specifically, we use the NLLB-200-1.3B-Distilled and mt5-large (1.2B parameter) models for our experiments and train NLLB models using Fairseq Ott et al. (2019) and mT5 with HuggingFace Transformers Wolf et al. (2020). When training in two stages, first with synthetic data and later with human errors, we keep the state of the learning rate scheduler, following the fine-tuning approach rather than retraining as defined by Grundkiewicz et al. (2019). See Appendices B.2 and B.3 for further details.
@@ 3.3 Generation
Prompt engineering. We perform iterative prompt engineering, analyzing intermediate qualitative results and updating the prompt. For instance, we initially started with a simple 2-shot prompt (temperature = 0.1) asking GPT-3.5 to add grammatical and spelling mistakes into the input text but noticed that some error types were missing. We then improved the prompt by specifying the missing error types, adding two more examples, and upping the temperature. Our final prompt uses four examples and a model temperature of 1.0. See Appendix A for the prompts. We randomly pick the examples from each language’s train set for few-shot prompting. When comparing the prompting between GPT-4-Turbo and GPT-3.5-Turbo, we use an identical random set of examples to ensure comparability.
@@
Probabilistic errors. We generate rule-based synthetic errors as done in prior work Grundkiewicz et al. (2019); Náplava and Straka (2019); Palma Gomez et al. (2023); Luhtaru et al. (2024) using the same method and also employing the Aspell speller§§§aspell.net for replacing subwords.
@@
| Llama + 1M prob + gold | 72.59 | 54.72 | 68.14 | 80.37 | 53.19 | 72.92 | 78.22 | 67.65 | 75.85 |
| Llama + 1M BT + gold | 73.85 | 57.83 | 69.97 | 82.03 | 53.41 | 74.09 | 79.08 | 68.66 | 76.75 |
@@
3.4 Evaluation
For Estonian, we evaluate our models using the Estonian learner language corpus (EstGEC-L2)¶¶¶github.com/tlu-dt-nlp/EstGEC-L2-Corpus/, alongside a modified version of the MaxMatch scorer∥∥∥github.com/TartuNLP/estgec/tree/main/
M2_scorer_est, following Luhtaru et al. (2024). The Estonian scorer also outputs recall per error category, accounting for both other errors within the word order error scope and not accounting for these. We report the ones that do consider other errors separately. For Ukrainian, our evaluation methodology aligns with that of the UNLP 2023 Shared Task Syvokon and Romanyshyn (2023), utilizing the CodaLab platform for submissions to a closed test set that uses the ERRANT scorer for evaluationBryant et al. (2017). We follow the GEC+Fluency track setting since it encompasses a wider range of challenging errors. For German, we use the test set from the Falko-Merlin (FM) corpus Boyd (2018) that several works have reported their scores on and the original MaxMatch scorer Dahlmeier and Ng (2012).
4 Results
In this section, we evaluate the performance of Llama-based models for GEC and AEG tasks. We then compare the AEG effectiveness between NLLB and mT5 models against Llama-based models to see if smaller, more efficient models can generate quality data. Separately, we assess AEG through prompting with GPT-3.5-turbo versus Llama models with trained error generation. Finally, we examine the quality of generated errors against human data and probabilistic reverse-speller errors.
@@ 4.1 Artificial Error Generation and Correction with Llama
We compare LLama-based language model (LM) fine-tuning error corrections across three configurations: (1) the baseline approach of training exclusively on human error GEC data, (2) the established related work approach of training on probabilistic reverse-speller AEG data and then continuing training with human error GEC data, and (3) our approach of training on back-translation style AEG data produced by fine-tuned Llama-based models first, followed by fine-tuning on human data.
@@
The resulting scores are compared in Table 2, along with previous state-of-the-art (SOTA) scores and results of GEC via 4-shot prompting of GPT-4/GPT-4-turbo. Results show that llama-based models, further enhanced through continued pre-training, exhibit strong correction capabilities across our study languages. Even without synthetic data, these models outperform current state-of-the-art (SOTA) methods in Estonian and Ukrainian error correction, and are not too far behind in German, trailing the best score by less than two points. However, it’s important to note the discrepancy in model sizes for a fair comparison; our 7B Llama model significantly exceeds the NLLB-200-1.3B-Distilled model Team et al. (2022) used for Estonian Luhtaru et al. (2024) and the mBART model Tang et al. (2021) for Ukrainian Bondarenko et al. (2023) in size. At the same time, it is smaller than the 13B mT5-xxl model used for German Rothe et al. (2021).
@@
Incorporating synthetic data as a preliminary step to fine-tuning significantly enhances performance across all languages and synthetic data types. Notably, our back-translation style synthetic data consistently delivers superior precision and recall compared to the probabilistic reverse-speller (or probabilistic) method. This approach results in a 2-2.4 point increase in the F score relative to solely using gold data for fine-tuning. Conversely, the gains from using probabilistic reverse-speller data are more modest, ranging from 0.6 to 1.5 points, highlighting the enhanced utility of our learned AEG errors.
@@
Our systems consistently outperform GPT-4 models regarding precision across all languages studied. However, GPT-4 models exhibit higher recall rates for Estonian and German. This discrepancy indicates that while our systems are more accurate in identifying correct instances, GPT-4 models better retrieve a broader range of relevant errors in these languages. On the other hand, the performance of GPT-4 models on the Ukrainian test set is notably lower compared to other methods and languages.

### v2

@@
This study explores enhancing grammatical error correction (GEC) through artificial error generation (AEG) using large language models (LLMs). Specifically, we fine-tune Llama 2-based LLMs for error generation and find that this approach yields synthetic errors akin to human errors. Next, we train GEC Llama models with the help of these artificial errors and outperform previous state-of-the-art error correction models, with gains ranging between 0.8 and 6 F0.5 points across all tested languages (German, Ukrainian, and Estonian). Moreover, we demonstrate that generating errors by fine-tuning smaller sequence-to-sequence models and prompting large commercial LLMs (GPT-3.5 and GPT-4) also results in synthetic errors beneficially affecting error generation models. We openly release trained models for error generation and correction and all the synthesized error datasets for the covered languages.
@@ 1 Introduction
The grammatical error correction (GEC) task aims to correct spelling and grammatical errors in text, making it valuable for a wide range of people. The best-performing GEC approaches currently use deep learning models (Junczys-Dowmunt et al., 2018; Omelianchuk et al., 2020; Rothe et al., 2021, and several others), which are known to be data-hungry. Simultaneously, the availability of openly accessible error correction data is severely limited, even for languages typically considered high-resource in other tasks, such as German and Arabic Bryant et al. (2023). This lack of data complicates the development of effective GEC systems for these and other even less-resourced languages.
@@
The scarcity of GEC data is commonly addressed through the creation of synthetic data, where errors are automatically added into correct sentences – also called artificial error generation (AEG). In low-resource settings, the overwhelmingly most employed approach for AEG is applying random probabilistic perturbation (deletion, insertion, replacement) of words and/or characters (Grundkiewicz et al., 2019; Rothe et al., 2021; Náplava and Straka, 2019, and others). Alternatives include usage of intricate hand-crafted rules and confusion sets Rozovskaya and Roth (2010); Xu et al. (2019); Kara et al. (2023); Bondarenko et al. (2023) and automatically learning to generate errors Xie et al. (2018); Kiyono et al. (2019); Stahlberg and Kumar (2021) – also referred to as back-translation (BT)***by analogy with the machine translation technique Sennrich et al. (2016). However, to the best of our knowledge, none of the related work on AEG uses pre-trained foundation models or applies this methodology in a low-resource setting.
@@
This gap is precisely the focus of the present work: we are using pre-trained language models for synthetic error generation and demonstrate the simplicity and effectiveness of the approach in low-resource scenarios. We approach the task by fine-tuning open large language models (LLMs) based on Llama 2 Touvron et al. (2023b) for error generation and correction, resulting in quality AEG data and state-of-the-art GEC models even when very limited human error data is available. Our analysis shows that the resulting errors can be categorized similarly to human errors. We also compare fine-tuning approach to prompting commercial LLMs (GPT-3.5 and GPT-4: OpenAI, 2023) to perform AEG, as well as include other open models commonly employed for GEC and tune them for AEG: mT5 Rothe et al. (2021); Palma Gomez et al. (2023) and NLLB Luhtaru et al. (2024).
@@
Our final goal and evaluation setting is improving GEC for languages with limited GEC data. In particular, we focus on German, Ukrainian, and Estonian GEC. When pre-trained on our LLM-generated synthetic errors, the resulting GEC models achieve the best current results on the included benchmarks in all three evaluated cases, including previous state-of-the-art and 4-shot GPT-4.
We publicly release AEG and GEC models from our work and the generated data. The datasets include one million sentences for German, Ukrainian, and Estonian, each processed with three different models, as well as an additional set of 100k sentences with GPT models.
@@
We show that pre-trained language models can be fine-tuned to generate high-quality synthetic errors even with limited data.
@@
We openly release GEC and AEG models as well as AEG datasets and implementation of training and inference to facilitate future research††† github.com/TartuNLP/gec-llm.
@@ 2 Related Work
The use of synthetic data is a common concept in GEC. The first effective neural method proposed by Junczys-Dowmunt et al. (2018) approaches GEC as low-resource Machine Translation (MT) translating from erroneous text to correct text, making it a relatively resource-heavy method encouraging synthetic data generation. Over the years, there have been different approaches to deliberately introducing errors into monolingual text, like rule-based and probabilistic methods, methods based on confusion sets and error patterns, models trained for error generation and using round-trip translation Bryant et al. (2023).
@@
Learned methods of error generation typically require more resources. Before the widespread adoption of transformers and MT, various studies explored alternative approaches for training models for error generation. For instance, Felice and Yuan (2014) and Rei et al. (2017) utilized statistical machine translation to generate errors, while Xie et al. (2018) and Yuan et al. (2019) experimented with CNNs for this purpose. Additionally, Kasewa et al. (2018) investigated using RNN-based sequence-to-sequence models with attention mechanisms.
Moving towards more modern MT architectures, Htut and Tetreault (2019) tested various model frameworks, including transformers, and Kiyono et al. (2019) specifically employed transformer models. Both of the latter studies trained models from scratch, utilizing datasets ranging from approximately 500,000 to over a million error correction examples to train the AEG system. In contrast, our work generates up to 1 million sentences with synthetic error while using between 9k and 33k human error sentences to fine-tune the base models.
@@
To the best of our knowledge, no works focus specifically on error generation by LLMs; however, several studies have evaluated the performance of commercial LLMs in this task. Fang et al. (2023b) found that while GPT-3.5 performs significantly worse than other systems in terms of precision, it excels in recall. Similar results were reported by Wu et al. (2023), who observed that GPT models tend to overcorrect rather than undercorrect errors. This finding is also supported by Coyne et al. (2023a), who noted that GPT models are particularly strong at making fluency edits. While there are few studies that use or evaluate open-source LLMs, Zhang et al. (2023) explore the use of the LLaMA model Touvron et al. (2023a) for writing assistance.
@@ 3 Methodology and Experiments
The primary target of our work is to apply generative language models to AEG via fine-tuning. Additionally, we experiment with prompting LLMs to perform the same task and include two seq2seq models that are fine-tuned to do the same.
The efficiency of proposed AEG solutions is evaluated by using them to improve GEC. Thus, we also fine-tune generative LLMs to perform the GEC task and compare the results to prompting-based GEC results and related work. The general pipeline of our approach is straight-forward:
@@
Fine-tune an LLM to generate errors using human error data, with correct sentences as input and sentences with errors as output.
@@
Apply that AEG LLM to correct sentences in order to add a synthetically erroneous counterpart.
@@
Fine-tune an LLM on that synthetic dataset to correct grammatical errors. Equivalent to Step 1, with the sentence pair direction reversed.
@@
Continue fine-tuning GEC LLM on the smaller dataset with human errors.
@@ 3.1 Data
We use two distinct types of data in our work. Firstly, we rely on datasets containing examples of corrections to train our error generation systems and correction models. Secondly, we incorporate monolingual data to create synthetic datasets by introducing errors. See an overview of used data in Table 1.
@@ 3.2 Models and Training
Llama-2-based models. We fine-tune models that have been enhanced with bilingual capabilities using continued pre-training from Llama-2-7B (Touvron et al., 2023b). For Estonian, we use Llammas-base‡‡‡huggingface.co/tartuNLP/Llammas-base Kuulmets et al. (2024), and for German, LeoLM§§§huggingface.co/LeoLM/leo-hessianai-7b. For Ukrainian, we apply continued pre-training to replicate the conditions of Estonian LLM by training with 5B tokens from CulturaX (Nguyen et al., 2023) with 25% of the documents being in English and the rest in Ukrainian following Kuulmets et al. (2024). For GEC and AEG fine-tuning, we formatted the training data with a prompt (see Table 12 and 13) loosely based on Alpaca (Taori et al., 2023). During fine-tuning, the loss is calculated on the tokens of the correct sentence. Fine-tuning details (including hyperparameters) are discussed in Appendix A.1.
Other models we use are NLLB Team et al. (2022) and mT5 Xue et al. (2021). Specifically, we use the NLLB-200-1.3B-Distilled and mt5-large (1.2B parameter) models for our experiments and train NLLB models using Fairseq Ott et al. (2019) and mT5 with HuggingFace Transformers Wolf et al. (2020). When training in two stages, first with synthetic data and later with human errors, we keep the state of the learning rate scheduler, following the fine-tuning approach rather than retraining as defined by Grundkiewicz et al. (2019). See Appendices A.2 and A.3 for further details.
@@ 3.3 Generation
Prompt engineering. We perform iterative prompt engineering, analyzing intermediate qualitative results and updating the prompt. For instance, we initially started with a simple 2-shot prompt (temperature = 0.1) asking GPT-3.5 to add grammatical and spelling mistakes into the input text but noticed that some error types were missing. We then improved the prompt by specifying the missing error types, adding two more examples, and upping the temperature. Our final prompt uses four examples and a model temperature of 1.0. See Appendix D for the prompts. We randomly pick the examples from each language’s train set for few-shot prompting. When comparing the prompting between GPT-4-Turbo and GPT-3.5-Turbo, we use an identical random set of examples to ensure comparability.
@@
Probabilistic errors. We generate rule-based synthetic errors as done in prior work Grundkiewicz et al. (2019); Náplava and Straka (2019); Palma Gomez et al. (2023); Luhtaru et al. (2024) using the same method and also employing the Aspell speller¶¶¶aspell.net for replacing subwords.
3.4 Automatic Evaluation of Models
For Estonian, we evaluate our models using the Estonian learner language corpus (EstGEC-L2)∥∥∥github.com/tlu-dt-nlp/EstGEC-L2-Corpus, alongside a modified version of the MaxMatch scorer******github.com/TartuNLP/estgec, following Luhtaru et al. (2024). The Estonian scorer also outputs recall per error category, accounting for both other errors within the word order error scope and not accounting for these. We report the ones that do consider other errors separately. For Ukrainian, our evaluation methodology aligns with that of the UNLP 2023 Shared Task Syvokon and Romanyshyn (2023), utilizing the CodaLab platform for submissions to a closed test set that uses the ERRANT scorer for evaluation Bryant et al. (2017). We follow the GEC+Fluency track setting since it encompasses a wider range of challenging errors. For German, we use the test set from the Falko-Merlin (FM) corpus Boyd (2018) that several works have reported their scores on and the original MaxMatch scorer Dahlmeier and Ng (2012).
3.5 Human Evaluation of Generated Data
In addition to evaluating the quality of our data in terms of its usefulness for training better models, we perform a detailed evaluation of generated data in Estonian. We apply the same annotation scheme Allkivi-Metsoja et al. (2022) used for annotating test and development sets to artificially generated sentences. This comparison allows us to assess the error distribution between the training data and generated data and to see whether the errors can be categorized into the same classes.
We select 100 random sentences from sets generated by Llama-based models, GPT-3.5-Turbo and GPT-4-Turbo††††††We also considered annotating probabilistic denoising errors, but these contained very few edits that could be categorized based on the annotation scheme., for annotation and also annotate 100 sentences from the training set. We add labels for problematic errors generated by the model, such as hallucinations and truncation of words important for understanding the meaning of sentence (HALL), synonym swaps (SYN), optional edits (O), corrections of mistakes in original sentences (INACC), and transformations that make the original word unrecognizable (UNREC).
4 Results
@@
| Llama + prob + gold | 72.59 | 54.72 | 68.14 | 80.37 | 53.19 | 72.92 | 78.22 | 67.65 | 75.85 |
| Llama + BT + gold | 73.85† | 57.83† | 69.97† | 82.03 | 53.41 | 74.09 | 79.08† | 68.66 | 76.75† |
@@
In this section, we evaluate the performance of Llama-based models for GEC and AEG tasks. We then compare the AEG effectiveness between NLLB and mT5 models against Llama-based models to see if smaller, more efficient models can generate quality data. Separately, we assess AEG through prompting with GPT-3.5-turbo versus Llama models with trained error generation. Finally, we examine the quality of generated errors against human data and probabilistic reverse-speller errors and compare the error type distributions for Estonian.
@@ 4.1 Artificial Error Generation and Correction with Llama
We compare LLama-based LLM fine-tuning error corrections across three configurations: (1) the baseline approach of training exclusively on human error GEC data, (2) the established related work approach of training on probabilistic reverse-speller AEG data and then continuing training with human error GEC data, and (3) our approach of training on back-translation style AEG data produced by fine-tuned Llama-based models first, followed by fine-tuning on human data.
@@
The resulting scores are compared in Table 2, along with previous state-of-the-art (SOTA) scores and results of GEC via 4-shot prompting of GPT-4/GPT-4-turbo. Results show that Llama-based models, further enhanced through continued pre-training, exhibit strong correction capabilities across languages in our study. Even without synthetic data, these models outperform current SOTA methods in Estonian and Ukrainian error correction, and are not too far behind in German, trailing the best score by two points. When comparing our 7B Llama model to others, there are significant differences in model sizes and data usage that need to be considered for a fair evaluation. Our 7B Llama model is substantially larger than the NLLB-200-1.3B-Distilled model used for Estonian Luhtaru et al. (2024) and the mBART model used for Ukrainian Bondarenko et al. (2023). However, it is smaller than the 13B gT5-xxl model, which represents the current state-of-the-art for German text-only data Rothe et al. (2021), while it is larger than the multimodal German model incorporating both text and speech data Fang et al. (2023a). In terms of synthetic data usage, our model is trained with one million sentences, which contrasts with the six million sentences per language used by Luhtaru et al. (2024) in their multilingual training approach, and the smaller, more carefully crafted synthetic datasets used by Bondarenko et al. (2023). Notably, all these models rely on the same human-labeled data, ensuring consistency in that aspect.
@@
Incorporating synthetic data as a preliminary step to fine-tuning significantly enhances performance across all languages and synthetic data types. Notably, our back-translation style synthetic data consistently delivers superior precision and recall compared to the probabilistic reverse-speller method. This approach results in a 2-2.4 point increase in the F0.5 score relative to solely using gold data for fine-tuning. Conversely, the gains from using probabilistic reverse-speller data are more modest, ranging from 0.6 to 1.5 points, highlighting the enhanced utility of our learned AEG errors.
@@
Our systems consistently outperform GPT-4 models in terms of precision across all languages studied. However, GPT-4 models exhibit higher recall rates for Estonian and German. This discrepancy indicates that while our systems are more accurate in identifying correct instances, GPT-4 models better retrieve a broader range of relevant errors in these languages.GPT-4’s performance on the Ukrainian test set is significantly lower compared to other methods and languages, likely due to the distinctive features of the dataset. Unlike the Estonian and German datasets, the Ukrainian set contains a higher proportion of punctuation errors (43%) and has a two times smaller error rate than German (8.2 vs 16.8) Syvokon et al. (2023). Since recent studies show that GPT models struggle with punctuation errors and tend to make more extensive changes to sentences Katinskaia and Yangarber (2024), this likely explains the variation in performance.